# VizAbility: Enhancing Chart Accessibility with LLM-based Conversational Interaction

### v2

@@ 4.3. Exploring Underlying Data Table via Keyboard
VizAbility offers users the flexibility to switch between the tree view and a conventional raw data table view (see options displayed in the buttons in Figure 1)— D5. While the tree view facilitates structured exploration based on visual encoding, the data table provides additional advantages like sorting features, enabling users to quickly access specific data values and patterns. Upon selecting a chart, users are presented with two options corresponding to the two modules: tree view and data tables. Clicking on either option activates the rendering of the chosen module on the page. Any user interactions specific to a module are retained during module switches. We disable navigation queries in the data table module as screen readers provide a slew of keyboard navigation shortcuts such as moving between headers and cells. The data table module supports the alphabetical/numeric sorting by each column.
@@ 4.4. Rapid Chart Probing via Conversational Interaction
The image illustrates a Q&A Pipeline that processes user queries through refinement and classification, leading to analytical, visual, contextual, and navigation queries, and integrates in-context learning to generate responses.
The use of keyboard navigation of chart content and its associated data can convey a clear picture of how a chart looks to blind users (Kim et al., 2023a), as well as intricate data details that constitute the chart. However, it can also be cumbersome to navigate individual encoding channels and data values or derive aggregate measures on the go. We incorporate the advanced capabilities of LLMs for question-answering. We outline the types of queries supported and explain the synergy between keyboard navigation and speech inputs to improve interaction with charts.
@@ 4.4.1. Data Set
We utilized a prior study’s data set, comprising 979 BLV user questions spanning four visual stimuli (bar, line, scatter, and map) for the development and quantitative evaluation of VizAbility. These questions were gathered through a wizard-of-oz study, where a human facilitator acted as a question-answering system. We reconstructed the visualization images into Vega-Lite specifications and partitioned the questions into analytical, visual, and contextual queries, which we derived from the question taxonomy in prior work (Kim et al., 2020, 2023b) (We explain their definitions in Section 4.4.2). We then partition the pool of questions once more into an 80/20 split between the testing and validation sets via stratified random sampling so that there is a proportionate representation of each query type amongst both sets.
The ground truths for the testing and validation sets were generated manually. Each user query within the data set has an accompanying ground truth classification, expressed as either analytical query, visual query, or contextual query, as well as a ground truth for the query response, for which we emphasized verboseness. For instance, the ground truth response to the question “What is the vaccination rate of South Africa” is “The vaccination rate for South Africa is 36%”, as opposed to the more concise “36%”. The presence of a ground truth enables us to evaluate both the quantitative and qualitative aspects of the response yielded by VizAbility.
Line Chart Bar Chart Scatterplot Choropleth Map Total
Analytical Query 117 137 196 155 605
Visual Query 36 21 37 32 126
Contextual Query 8 8 21 9 46
Navigation Query N/A N/A N/A N/A N/A
Total 161 166 254 196 777
Table 1. Distribution of test data for four types of queries across different chart formats, including line chart, bar chart, scatterplot, and choropleth map, indicating the prevalence of analytical queries in the dataset.
@@ 4.4.2. Supported Query Types
As described above, we derived the following query types from the question set from BLV individuals (Kim et al., 2023b). These query types have been discussed in existing chart question answering work (Kim et al., 2023b, 2020; Masry et al., 2022), while we consolidate similar queries into three query types: data-driven queries, visual queries, and context-seeking queries. We created a new navigation query type to accommodate natural language questions to control keyboard navigation.
As previously described, we have derived four query types from the question sets of BLV individuals(Kim et al., 2023b). These types have also been discussed in existing chart question answering research(Kim et al., 2023b, 2020; Masry et al., 2022). In our work, we have consolidated similar queries into three distinct types: data-driven queries, visual queries, and context-seeking queries. Furthermore, to better accommodate natural language questions for controlling keyboard navigation, we have introduced a new category, the navigation query type.
@@
Navigation queriesD4 are a category we introduced to enhance user experience. These queries are tailored to the synergy between keyboard navigation and conversational interaction. For instance, to reduce cumbersome keyboard navigation and assist users in orientation, questions such as “How can I get to the X-axis” (direction) or “Where am I?” (orientation) can be beneficial. Our motivation for this stems from a previous empirical study (Kim et al., 2023a), where blind individuals highlighted such challenges with Olli’s tree view.
@@ 4.4.3. Query Classification
On the other hand, navigation queries are a novel consideration and are therefore absent from the validation set. Thus, we derived example navigation queries from our formative evaluation with a blind design partner. Whereas analytical, visual, and contextual queries entail a broad range of questions, the general format of a navigation question is quite predictable, given its specific intention to either orient the user or direct the user to another node within the tree view. Question syntax—as opposed to content—therefore proves to be the only variable that the LLM must account for; e.g., Take me to the y-axis..
@@ 4.4.4. Query-Specific Prompting
Once the starting and ending points have been identified, we employ a breadth-search algorithm that returns string instructions of the shortest path, which users can then manually follow at their own discretion. To avoid repetition we coalesce instructions whenever possible. For instance, for instructions pertaining to the navigation between two sibling nodes, we convert “Press the right arrow key. Press the right arrow key. Press the right arrow key.” to “Press the right arrow key 3 times.” We opted for this approach as opposed to automatically moving the user to their desired ending point with the rationale that autonomy and transparency are crucial for our intended audience.
@@ 5. Evaluation: Q&A Performance Benchmark
Figure 7. Quantitative results display the distributions for classification accuracy and quality ratings (via a 5-point likert scale) for response to user questions in the testing set.
Figure 7. Quantitative results display the distributions for classification accuracy and quality ratings (via a 5-point likert scale) for response to user questions in the testing set.
@@ 5.1. Classification Evaluation
We compared the classification result of VizAbility to the ground truth query type. We used a relaxed comparison, allowing for any potential discrepancies in formatting (such as the addition of a white space or newline character by the LLM). If there is a 1:1 correspondence, we output “Correct Classification”. Otherwise, we output “Incorrect Classification”.
The evaluation of our testing set yielded a classification accuracy of 88.5% ($688777\frac${688}{777}divide start_ARG 688 end_ARG start_ARG 777 end_ARG). The $8888$ user queries which were incorrectly classified by the LLM consisted of $5252$ queries that could not be classified (signifying overtly vague or impossible to answer questions) and $3636$ queries that were classified into a query type that did not correspond with the ground truth. For instance, VizAbility could not properly classify the following query directed at the line chart, which depicts the plummeting inventory of homes for sale and makes no mention of geographic data: “Where are the homes sold from? Country, city or state what does it relate to?”.
@@ 5.2. Question Response Evaluation
5.2.1. Evaluation Pipeline
We employed GPT-4 to evaluate the quality of natural language responses, which is an established practice in natural language evaluation (Liu et al., 2023; Wang et al., 2023; Fu et al., 2023). With us identifying trustworthiness and transparency as crucial factors, we wanted to reflect this fact by emphasizing explanatory responses over more concise ones. We adopted a Likert scale evaluation prompt, inspired by Liu et al. (Liu et al., 2023), which framed a response’s ‘correctness’ in terms of its coherence with the ground truth. This coherence metric ranged from 1-5, but for clarity, we adapted it to the Likert Scale of Quality: [Very Poor, Poor, Fair, Good, Very Good]. Refer to Table 2 to view how we defined each of the increments on the Likert Scale of Quality.
@@ 5.2.1. Evaluation Pipeline
The evaluation prompt presented two responses: Response A and Response B, with Response A acting as the ground truth. GPT4 was directed to assess the coherence of Response B in relation to Response A. We refrained from revealing which response was the ground truth or our own creation, as we deemed this indication to be extraneous information for the LLM, given its task to solely assess the coherence of two responses. Given that the coherence metric acts as an umbrella term for response evaluation (encompassing the metrics of correctness, phrasing, verboseness, etc.) we further prime the LLM through few-shot prompting. We populate the prompt with a manually selected sample of questions, VizAbility responses, and their corresponding ground truth that we derive from the validation set. To further contextualize LLM evaluation responses, we direct the Language Model to append a one-sentence rationale to its Likert Score. During prompting, if the score deviated from the 1-5 range, GPT4 reassessed its evaluation of Response B. The results were formatted as “Score: $\{${coherence score}}\}} Rationale $\{${rationale behind coherence score}}\}}”.
5.2.2. Overall Response Evaluation Results
@@
[…]
5.2.3. Partitioning evaluation results based on user query classification
[…]
5.2.4. Partitioning evaluation results based on the user query’s relevance to the data set
User queries such as “What was the interest rate in 2015?” do not pertain at all to their corresponding chart, which in this instance is the line chart depicting the decreasing inventory of homes for sale over the years. This observation motivated us to partition the final evaluation results on the basis of the question’s relevance to the chart. Our 3-point scale for relevance spanned the values [Low, Medium, High]. Low relevance was defined as bearing no relevance whatsoever to the chart. Questions of medium relevance are on topic but have some ambiguities such as unclear labels or missing information. Finally, questions of high relevance are concrete, concise, and aligned with the chart’s contents.
[…]
5.2.5. Partitioning evaluation results by partitioning based on the response leniency and query broadness
User queries exhibit diverse natures, reflected in our query classification. Notably, while both ”What is the vaccination rate of Columbia?” and ”What is the average vaccination rate in South America?” are classified as analytical, the former refers directly to a data set variable, whereas the latter involves a computed value prone to formatting variations.
Hence, our evaluation of VizAbility should take into account the implied broadness of a user query. This leniency extends to ”Why” queries, which allow for even more diverse and varied responses than their “What” counterparts. Given this observation, we developed a new dimension for our quantitative evaluation: the “Afforded Leniency” metric. Questions that involve the retrieval of variables from the dataset are afforded “Low” leniency; those that require computations or generation of lists are manually classified as “Medium” afforded leniency. Finally, questions that require detailed, verbose responses– such as “What does the data represent?”– are afforded “High” leniency.
[…]
5.3. Comparisons to an existing system
[…]

### v3

@@ 4.3. Exploring Underlying Data Table via Keyboard
VizAbility offers users the flexibility to switch between the tree view and a conventional raw data table view (see options displayed in the buttons in Figure 1)— D5. While the tree view facilitates structured exploration based on visual encoding, the data table provides additional advantages like sorting features, enabling users to quickly access specific data values and patterns. We disable navigation queries in the data table module as screen readers provide a slew of keyboard navigation shortcuts such as moving between headers and cells. The data table module supports the alphabetical/numeric sorting by each column.
@@ 4.4. Rapid Chart Probing via Conversational Interaction
[…]
Keyboard navigation helps blind users grasp the chart’s visual structure and data, but can be cumbersome for higher-level comprehension, such as computing aggregates or understanding overall visuals. We integrate LLMs for intuitive chart question-answering and enhance interaction by combining keyboard navigation with speech inputs.
Query Type Line Chart Bar Chart Scatterplot Choropleth Map Total
Analytical Query 147 168 242 193 750
Visual Query 42 28 54 41 165
Contextual Query 16 14 28 12 70
Navigation Query 12 12 12 12 48
Total 217 222 336 258 1033
Table 1. Distribution of data for four types of queries across different chart formats, including line chart, bar chart, scatterplot, and choropleth map, indicating the prevalence of analytical queries in the dataset.
@@ 4.4.1. Data Set
We utilized a prior study’s data set (Kim et al., 2023b), comprising 979 BLV user questions spanning four visual stimuli (bar, line, scatter, and map) for the development and quantitative evaluation of VizAbility (Table 1). We then partition the pool of questions once more into an 80/20 split between the testing and validation sets via stratified random sampling so that there is a proportionate representation of each query type amongst both sets.
The questions were gathered through a wizard-of-oz study, where a human facilitator acted as a question-answering system. We reconstructed the visualization images into Vega-Lite specifications and partitioned the questions into analytical, visual, and contextual queries, which we derived from the question taxonomy in prior work (Kim et al., 2020, 2023b). In addition, we employed GPT4 to generate 12 example navigation queries for each of the four visual stimuli, appending them to the prior data set of 979 BLV user questions. We define each query type in Section 4.4.2.
How did we derive the question taxonomy?
The original taxonomy by Kim et al.(Kim et al., 2023b) offers more fine-grained categories, such as granular analytical tasks like retrieving values and finding extremum(Amar et al., 2005). They also categorized queries as visual vs. non-visual and look-up vs. compositional, similar to previous chart QA systems (Kim et al., 2020), while separating non-data queries. In our work, we reclassify these queries, as the detailed categorization is not needed for users and may not enhance the system’s performance, considering the advanced language comprehension capabilities of LLMs. First, we consolidated data-driven and visual tasks into analytical (or non-visual) and visual queries, respectively. We do not separately consider look-up vs. compositional, as we expect LLMs to handle them without differentiation. Finally, we consolidated non-data queries into contextual queries.
What additional metrics did we introduce?
We also elaborate on the previous taxonomy by introducing two new binary metrics: open-endedness and answerability. Open-ended queries allow for multiple valid interpretations and responses, and often invoke questions like “Why?”. Open-ended questions may also involve computations, especially when the operation parameters are ambiguous: “And what was that temperature?”. Answerable queries are relevant to the data set. By contrast, unanswerable queries are irrelevant and cannot be feasibly addressed using the corresponding chart.
How did we generate navigation queries?
We used GPT-4 with few-shot prompting to generate navigation queries with an equal distribution of orientation and wayfinding queries. To simulate real user chart interaction, we include the corresponding active element in the treeview for each query.
How did we generate ground-truths?
The ground truths for the testing and validation sets were manually generated by two researchers independently and then merged by resolving conflicts. The process involved reading charts, calculating numbers, and searching for information online. Throughout the generation process, we emphasized verboseness. For instance, the ground truth response to the question “What is the vaccination rate of South Africa” is “The vaccination rate for South Africa is 36%”, as opposed to the more concise “36%”. To prioritize the model’s ability to translate ambiguous user wording into precise nodes within the treeview, the ground truth responses for navigation queries consist solely of starting and ending nodes.
@@ 4.4.2. Supported Query Types
Navigation queriesD4 are a category we introduced to enhance user experience. These queries are tailored to the synergy between keyboard navigation and conversational interaction. For instance, to reduce cumbersome keyboard navigation and assist users in orientation, questions such as “How can I get to the X-axis” (wayfinding) or “Where am I?” (orientation) can be beneficial. Our motivation for this stems from a previous empirical study (Kim et al., 2023a), where blind individuals highlighted such challenges with Olli’s tree view.
@@ 4.4.4. Query-Specific Prompting
Once the starting and ending points have been identified, we employ a breadth-search algorithm that returns string instructions of the shortest path, which users can then manually follow at their own discretion. To avoid repetition we coalesce instructions whenever possible. For instance, for instructions pertaining to the navigation between two sibling nodes, we convert “Press the right arrow key. Press the right arrow key. Press the right arrow key.” to “Press the right arrow key 3 times.” We initially opted for this approach as opposed to automatically moving the user to their desired ending point with the rationale that autonomy and transparency are crucial for our intended audience.
@@ 5.1. Classification Evaluation
Our evaluation yielded an overall classification accuracy of 87.39%. Table 2 presents detailed results, including precision, recall, and F1-scores for each class. Table 3 provides examples of misclassified instances.
Query Type Precision Recall F1 Count
Analytical 90.96% 93.10% 92.02% 551
Contextual 64.65% 67.37% 65.98% 95
Navigation 100% 97.50% 98.73% 40
Visual 89.09% 74.81% 81.33% 131
Table 2. Quantitative results contextualize VizAbility’s classification accuracy of 87.39% through the lenses of Precision, Recall, and F1 scores. The system is most proficient at classifying analytical and navigation queries, as indicated by the significantly higher F1 Scores attributed to these query types.
True Class Predicted Class Example Text
Analytical Contextual Does the data include booster shot?
Analytical Visual Is Asia in the medium upper range in 60 to 80 area?
Contextual Analytical What is the percentage at which vaccinated population reach herd immunity?
Contextual Analytical What is the source of this data?
Contextual Visual Describe a scatterplot.
Contextual Visual Can I get a map of the US?
Visual Analytical How many years in total are represented on the X axis?
Visual Analytical Which part of Asia has a darker shade of green…
Visual Analytical What is the trend for Brown?
Visual Contextual Is there anything specific you want me to look for in this chart?
Navigation Visual …am I closer to the start or the end of the Y-axis?
Table 3. Additional Examples of Misclassification Cases
The model was effective in identifying analytical queries, with a 93.10% ($\frac{513}{550}$) recall and 90.96% ($\frac{513}{564}$) precision. Analytical queries often contain distinct computationally-heavy language, such as “correlation” and “average decrease,” as seen in examples like “What’s the correlation between GDP per capita and population?” and “For the time from September 2019 to September 2020, what was the average decrease in homes for sale?” This specific language may aid in differentiating these queries from broader contextual ones, potentially leading to a higher overall accuracy.
Of the $95$ queries designated as contextual in the ground truth, 67.37% (64) were classified as such by VizAbility. In addition to the 64 true positives, the model incorrectly classified 30 analytical queries and 5 visual queries as contextual in nature - contributing to a precision of 64.65% ($\frac{64}{99}$). Poorly classified contextual queries like “What is the source of this data?” and “Can I get a map of the US?” may indicate that the model often emphasized individual words over the overall meaning of queries. For instance, the word “data” likely led the model to classify the first query as analytical. Similarly, the word “map” likely prompted the model to mistakenly identify the second query as visual.
[…]
VizAbility’s proficiency in distinguishing navigation queries may be attributed to the distinct and uniform structure to which most GPT4-generated example queries conform; i.e., queries either assume the role of wayfinding or orientation questions. The results, showing a 100% ($\frac{39}{39}$) precision and 97.50% ($\frac{39}{40}$) recall, suggest that VizAbility is effective in distinguishing these tasks from the typical data retrieval or lookup associated with analytical and visual queries. Nonetheless, the misclassified query “While exploring inventory values between 800000 and 1000000, am I closer to the start or the end of the Y-axis?” hints that the model might sometimes give undue weight to individual words. For example, the reference to ‘the Y-axis’, a visual element of the chart, could have led the model to categorize the navigation query as visual.
Figure 7. Quantitative results display the distributions of quality ratings (via a 5-point likert scale) for VizAbility responses. For more granularity, the results are also partitioned by query type, chart type, and question characteristics. VizAbility’s performance is considerably higher than the GPT-4V baseline.
[…]
@@ 5.2. Question Response Evaluation
Our evaluation of the response quality is twofold. First, two researchers manually inspected the quality of the responses. Based on this ground-truth evaluation, we established an LLM-based evaluation using GPT-4, which is intended to facilitate scalable, automatic evaluation for efficient benchmarking tests, following recent literature in natural language evaluation (Liu et al., 2023; Wang et al., 2023; Fu et al., 2023; Li et al., 2024). We describe each of these procedures below.
5.2.1. Human Evaluation
We evaluated each pair of system response and corresponding ground truth individually based on a five-point Likert scale: [Very Poor, Poor, Fair, Good, Very Good]. Our focus was on a response’s ‘correctness’ in terms of its coherence and consistency with the ground truth. Our assessment scheme favored explanatory responses over overly brief ones, in line with our recognition of trustworthiness and transparency as essential factors. Table 4 shows how we defined each increment on the Likert Scale, which is further elaborated in the GPT4-based assessment in Section 5.2.2.
@@ 5.2.1. Human Evaluation
Of the $817$ user queries, $632$ or $77.36\%$ were deemed “Very Good” in the human assessment. Responses rated as “Very Good” often restated the user’s question, formatted quantitative data correctly, and included contextual information. For example, in response to the query “What continent has the highest vaccination rate?” related to a choropleth map, VizAbility answered, “The continent with the highest vaccination rate based on the percentage of fully vaccinated people in each country is South America, with an average percentage of 69.7%.” This response, more detailed than the ground truth “The continent with the highest vaccination rate is: South America”, demonstrates VizAbility’s ability to provide comprehensive answers. The distribution for Good, Fair, and Poor responses was $5.14\%$ or $\frac{42}{817}$, $4.65\%$ or $\frac{38}{817}$, and $7.71\%$ or $\frac{63}{817}$ respectively. The human assessment yielded $42$ or $5.14\%$ of questions as being ”Very Poor” in coherence to the ground truth. We elaborate on these findings and investigate relationships between the types of user queries and VizAbility’ responses below (Figure 7).
Partitioning based on query classification
VizAbility exhibited the greatest accuracy in answering navigation questions. 87.5% ($\frac{35}{40}$) of navigation queries received a “Very Good” assessment, which equates to the correct translation of the user query into concrete starting and end nodes within the treeview. For instance, VizAbility correctly parsed the question “What’s the quickest path to get from the top of the tree to inventory values above 1400000?” and identified “Starting Point: ‘A line chart. With axes Date and Number of Homes for Sale’; Starting Address: 1; Ending Point: ‘Inventory is between 1400000 and 1600000’; Ending Address: 1.2.6”, which corresponds with the ground truth. User references to the treeview may not always be explicit (“top of the tree” $\rightarrow$ “a line chart…”). 82.11% ($\frac{78}{95}$) of responses to contextual queries were identified as being “Very Good”. This metric was slightly lower for both visual and analytical queries, from which 77.86% ($\frac{102}{131}$) and 75.68% ($\frac{417}{551}$) of responses had warranted a “Very Good” assessment respectively. Nevertheless, the distribution of assessment scores is relatively consistent across query types, with a tendency to skew towards “Very Good” (Figure 7).
Partitioning based on chart types
The distribution of assessment scores is similar across chart stimuli, suggesting that chart type may not have as significant an influence on overall system performance compared to other factors. Questions pertaining to the choropleth map had the highest frequency of yielding “Very Good” responses (79.61% or $\frac{164}{206}$). This is followed by responses for the scatter plot (77.65% or $\frac{205}{264}$), bar graph (77.27% or $\frac{136}{176}$), and line chart (74.27% or $\frac{127}{171}$). The low variability amongst chart stimuli highlights VizAbility’s versatility in addressing a wide range of data visualizations.
Partitioning on the answerability of queries
We found that of all the questions for which VizAbility’s response was prescribed a “Very Poor” rating, almost half ($45.24\%$ or $\frac{19}{42}$) had also been identified as being unanswerable (Figure 7). More notably, 76% ($\frac{48}{63}$) of “Poor” responses also happened to be unanswerable. For instance, referring to the choropleth map which depicts the percentage of the global population receiving at least one dose of the vaccine, the query “Which country has the fewest vaccinations?” is unanswerable because of the absence of precise counts of vaccines from the data set. Nonetheless, VizAbility fails to acknowledge that relative percentages do not equate to precise counts and generates the “Very Poor” output “Among the listed countries, Burundi has the fewest vaccinations.”. Instances like these skew the overall accuracy of the system towards “Very Poor”. By isolating just the answerable queries and recomputing accuracies, we find that 87.03% ($\frac{584}{671}$) of responses are now “Very Good” and consistent with the ground truth, increasing the mean assessment score from $4.42$ to $4.70$.
Partitioning based on the open-endedness of queries
70.10% ($\frac{143}{204}$) of system responses to open-ended queries were prescribed a score of “Very Good”. For the assessment of open-ended contextual queries, we emphasized VizAbility’s ability to not only draw on external information, but also support claims with relevant data derived from the chart. The contextual query, “Is there any other places that I can get some new information like a graph which shows by each region of the US?” with ground truth: “Yes, charts depicting inventory of houses for sale by region can be found on the internet.”, received the following “Very Good” response:
“Yes, you can find a similar graph that displays home sales data by different regions of the US. The existing dataset provides a line chart showing the number of homes for sale over time, and this type of data is often collected and analyzed at both the national and regional levels. Organizations such as the National Association of Realtors, Zillow, and Redfin regularly publish reports […]”
In this response, VizAbility not only explicitly identifies credible sources that the user can consult, but also clarifies those sources’ relevance to the line chart. Nonetheless, VizAbility proved to be more proficient in answering questions that were not open-ended. Of the $613$ user questions that are narrower in scope (not open-ended), $79.77\%$ yielded “Very Good” system responses - a value almost 10% higher than the reported accuracy for open-ended queries (Figure 7). This performance increase may be due to the narrow scope of non-open-ended questions, which could reduce the chances of computational or logical errors; e.g., scatter plot questions such as “What color is North America?” are straightforward. Consequently, VizAbility’s response, “The color that represents North America in the dataset is red.” is almost verbatim to the ground truth, “The color of North America on the scatter plot is: Red”. Similarly, responses to non-open-ended queries were also less frequently assessed as “Very Poor”; 3.92% vs. the rate of 8.82% for open-ended queries.
5.2.2. Automatic Evaluation using GPT4
We leverage LLMs for holistic evaluations that align better with human judgments compared to traditional metrics like BLEU, ROUGE, and BERTScore, which focus on simple text-level differences (Li et al., 2024).
Our evaluation prompt, inspired by Liu et al. (Liu et al., 2023), presented two responses to GPT4 (gpt-4-0125-preview): Response A and Response B, with Response A acting as the ground truth. GPT4 was directed to assess the coherence of Response B in relation to Response A. We refrained from revealing which response was the ground truth or our own creation, as we hypothesized this indication to be extraneous information for the LLM, given its task to solely assess the coherence of two responses. The lower Kendall $\tau$ score ($=0.2900$) with human ground-truths for a similar prompt during our iterative testing, which explicitly defined Response A as the ground truth, supported this decision.
@@ 5.2.2. Automatic Evaluation using GPT4
Our automatic evaluation prompt yielded a Kendall’s $\tau$ score of 0.5526 ($p<0.001$), signifying a strong correlation ($|\tau|\in[0.3,1.0]$) between the automatic assessments and human evaluation (Chiang and Lee, 2023). The exact distribution of score ratings for GPT4 assessments is as follows: “Very Poor”: 16.52% or $\frac{132}{817}$, “Poor”: 5.02% or $\frac{41}{817}$, “Fair”: 3.18% or $\frac{26}{817}$, “Good”: 12.85% or $\frac{105}{817}$, “Very Good”: 62.42% or $\frac{510}{817}$. GPT4 demonstrated more variability in its assessments, with fewer outputs clustered around the “Very Good” rating. More precisely, we observed a reduction of 122 “Very Good” assessments and an increase of 63 “Good” assessments between the human and LLM evaluations, which might signify that GPT4 had adopted stricter and more rigourous criteria. For the choropleth map query, “What is the highest vaccination rate in Africa?”, which yielded the system response, “The country with the highest vaccination rate in Africa based on the geographic map dataset is Rwanda.”, the human evaluation designated this response as “Very Good” in overall quality due to its clarity, explanatory language, and close correspondence with the ground truth: “The highest vaccination rate in Africa is 78.00% and belongs to Rwanda.”. Nonetheless, GPT4 attributed a score of “Good”, citing “Both Response A and Response B identify Rwanda as the country with the highest vaccination rate in Africa, although Response B does not provide the specific percentage rate.”
5.3. Baseline Comparisons
5.3.1. Comparison to a question-answering system with symbolic inputs
We compared our system to a similar system that focuses on chart reasoning. (Kim et al., 2020). Directly running the system proved challenging due to compatibility issues with outdated dependencies. A similar evaluation, using our dataset from blind participants, was conducted in prior work, revealing an overall factual accuracy rate of 16% (Kim et al., 2023b). This evaluation, however, was limited to just 245 queries from the total dataset. It excluded contextual queries and others deemed unanswerable due to ambiguous wording. Furthermore, the evaluation focused solely on questions related to bar and line charts, aligning with the system’s supported question types.
Overall, the system demonstrated some level of proficiency in answering straightforward value retrieval and extrema questions. The relatively low performance was mainly attributed to query comprehension and handling of diverse task types (e.g., yes/no and range questions). Seeking to maintain consistency with the prior system, we extracted data solely from the bar and line charts for a more fitting comparison. When narrowing the scope to these two types of visual stimuli, VizAbility reports $\approx 76.01\%$ accuracy for the line chart and $\approx 81.82\%$ for the bar chart, only considering ‘Very Good’ and ‘Good’ responses, signifying a significant improvement in user query handling.
5.3.2. Comparison to GPT-4V(ision) with image inputs
We conducted a comparative analysis between VizAbility and GPT-4 with Vision. Supplying GPT-4V with input images of the line chart, bar graph, scatter plot, and choropleth map, we generated responses to all $777$ analytical, contextual, and visual queries in the test set. We assessed the quality of GPT-4V responses using our GPT4-based automatic evaluation pipeline, which has been verified to correlate strongly with human evaluations. We omitted navigation queries because it is infeasible to provide just one reference image for the keyboard-navigable treeview. On the other hand, supplying the treeview text is essentially equivalent to using VizAbility.
The GPT4 automatic evaluation yielded the following distribution of scores: 29.86% or $\frac{232}{777}$ of responses received a “Very Good” assessment. 10.55% or $\frac{82}{777}$ of responses were “Good”, whereas 12.23% or $\frac{95}{777}$ were “Fair”. “Poor” and “Very Poor” responses occurred at equal frequencies: 23.68% or $\frac{184}{777}$. This discrepancy in performance between GPT-4V and VizAbility, which outputs “Very Good” responses at a frequency of 77.36%, can be attributed to GPT-4V lacking access to the raw data required for queries that require symbolic computations, resulting in inconclusive responses for certain prompts, such as “What is the average vaccination rate for countries in the world?”:
“The image you’ve provided is a world map that shows the share of the population receiving at least one dose of a COVID-19 vaccine… To determine the average vaccination rate for countries around the world, one would need to access the raw data from a reliable source such as the World Health Organization, a government health department, or a global health tracking website…”.
Furthermore, the responses outputted by GPT-4V tend to be excessively verbose. For instance, in the above response, GPT-4V initially strays away from the user question by providing an unnecessary and potentially redundant description of the choropleth map. On average, GPT-4V responses were longer by 97.69 words, which can detract from user readability and concision. Yet, GPT-4V often performed reasonably well on certain visual questions such as asking overall shapes of trends (e.g., “How is the graph moving?” “Would it be […] going up like hills and valleys type of trend?”).